Integrate with respect to x: ∫ 3e^x dx

a. e^x+c

b. 3x+c

c. 3e^x+c

d. 3e^x
1
Expert's answer
2013-05-23T08:55:41-0400
At first we can rewrite the given integral as follows:
∫ 3e^x dx=3 ∫ e^x dx
Using table of integration we obtain that ∫ e^x dx=e^x+const then 3 ∫ e^xdx=3*( e^x+const )=3e^x+c
So the right answer is c. 3e^x+c